OpenStudy (anonymous):

A cylinder has a radius that is decreasing at a rate of 2 cm per second while its height Is increasing at a rate of 3 cm per second . At what rate is the volume changing when the radius is 3 cm and the height is 2cm ?

OpenStudy (phi):

can you write down the formula for the volume of a cylinder?

OpenStudy (anonymous):

V = π r 2 h

OpenStudy (anonymous):

v=pir^2h

OpenStudy (phi):

can you take the derivative (with respect to time "t") of both sides of that equation?

OpenStudy (anonymous):

dv/dt=pi2rh*dh/dt

OpenStudy (phi):

ok on dv/dt on the right side (ignore the pi for the moment) we have to use the product rule \[ \frac{d}{dt} r^2 h = r^2 \frac{d}{dt} + h \frac{d}{dt} r^2\] can you finish that ?

OpenStudy (phi):

**\[ \frac{d}{dt} r^2 h = r^2 \frac{dh}{dt} + h \frac{d}{dt} r^2 \]

OpenStudy (phi):

the product rule is d( u v) = u dv + v du

OpenStudy (phi):

now use the power rule on r^2 and the chain rule d r^2 = 2 r dr

OpenStudy (phi):

d is short for \(\frac{d}{dt} \)

OpenStudy (anonymous):

dv/dt=pi*2r*dr/dh+r^2dh/dt

OpenStudy (phi):

ok except pi multiplies both terms on the right side

OpenStudy (phi):

\[ \frac{dv}{dt}= \pi \left( \frac{d}{dt} r^2 h \right) \]

OpenStudy (phi):

last step is replace the variables with the numbers given in the problem

OpenStudy (phi):

btw, you have dr/dh. that should read dr/dt so you should get dv/dt=pi* ( 2r*dr/dt+r^2 dh/dt)

OpenStudy (anonymous):

dv/dt=15pi cm/s?

OpenStudy (phi):

so you should get dv/dt=pi* ( 2r*h*dr/dt+r^2 dh/dt) we should use the equation editor. Keep leaving out terms. It should read \[ \frac{dv}{dt}= \pi \left( 2rh \dot{r} + r^2 \dot{h} \right) \] r=3 h=2 \(\dot{r} = 2\) \( \dot{h} = 3 \)

OpenStudy (phi):

Newton used \(\dot{x} = \frac{dx}{dt} \)

OpenStudy (anonymous):

coool

OpenStudy (phi):

It's not 15 pi

OpenStudy (anonymous):

ohhhh I left out a h

OpenStudy (anonymous):

3 pi

OpenStudy (anonymous):

I thought dr/dt would be -2cm/s since its decreasing at a rate of 2 cm/s

OpenStudy (phi):

yes. good catch.

OpenStudy (anonymous):

alright . thank you ;)

OpenStudy (phi):

yw